OpenStudy (anonymous):

While I can see that this equation must approach 1 as x approaches infinity. How would you go about notating and using limits to say so? y[x] = E^(rx) E^(bk) / (E^(rx) E^(bk)-1)

OpenStudy (anonymous):

\[y[x] = \frac{E^{rx} E^{bk} }{ (E^{rx} E^{bk}-1)}\]

OpenStudy (anonymous):

what are the values of \(E\) ?

OpenStudy (anonymous):

sorry.. E is the natural euler exponent ,

OpenStudy (anonymous):

is that supposed to be lower case?

OpenStudy (anonymous):

ok, and how about \(r\)? Is it a positive number?

OpenStudy (anonymous):

r is positive, and I think we can say b is too

OpenStudy (anonymous):

lets say they all are.. k too

OpenStudy (anonymous):

then\[\lim_{x \to \infty} \frac{e^{rx} e^{bk} }{ e^{rx} e^{bk}-1}=\lim_{x \to \infty} \frac{e^{rx} e^{bk} }{ e^{rx} e^{bk}}=1\]you can neglect \(1\) in the denom, because \(e^{rx} e^{bk}>>1\)

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

looks good to me :) so the constant 1, can be eliminated on grounds that the first terms are going to be insanely greater ... is there a need to cancel anything from that position?

OpenStudy (anonymous):

emm, no, It's fine from there

OpenStudy (anonymous):

so is this a general rule that constants are considered 0?

OpenStudy (anonymous):

yeah, you can use that fact

OpenStudy (anonymous):

sweet thank you.

OpenStudy (anonymous):

np

OpenStudy (michele_laino):

here, we can divide both numerator and denominator, by e^(rx), so we get: \[\Large y\left( x \right) = \frac{{{e^{rx}}{e^{bk}}}}{{{e^{rx}}{e^{bk}} - 1}} = \frac{{{e^{bk}}}}{{{e^{bk}} - {e^{ - rx}}}}\] Now, if r>0, then we have: \[\Large \mathop {\lim }\limits_{x \to + \infty } {e^{ - rx}} = 0\] then the requested limit value is: \[\Large \mathop {\lim }\limits_{x \to + \infty } y\left( x \right) = \mathop {\lim }\limits_{x \to + \infty } \frac{{{e^{bk}}}}{{{e^{bk}} - {e^{ - rx}}}} = \frac{{{e^{bk}}}}{{{e^{bk}}}} = 1\]

OpenStudy (michele_laino):

provided that: \[\Large r > 0\]

OpenStudy (anonymous):

thank you michele .. that's exactly what I was looking for.

OpenStudy (anonymous):

I knew there was some kind of cancellation I could use, but I couldn't see it.

OpenStudy (michele_laino):

:)